If (x,y,z)=3x²y-yz?, find grad at point (2,2,-1)


1
Expert's answer
2021-10-10T15:58:13-0400

By definition, a gradient of a function "f" is a vector function given by "\nabla f (x,y,z) = (\frac{\partial f}{\partial x}, \frac{\partial f}{\partial y}, \frac{\partial f}{\partial z})". Let us calculate each of partial derivatives :

"\frac{\partial f}{\partial x} = 6xy, \: \frac{\partial f}{\partial y}=3x^2-z, \: \frac{\partial f}{\partial z}=-y"

Putting "(x,y,z)=(2,2,-1)" gives us a vector "(24, 13, -2)".